# Towards Generalist Robot Learning from Internet Video: A Survey

### v3

@@ 6.3 Overcoming LfV Challenges
The missing action label problem in LfV can be partially addressed using alternative-action representations (see Section 4.2.1). We recommend future research to: (1) Scale existing methods to realistic and diverse internet video. (2) Explicitly compare the pros and cons of different action representations, and measure the extent to which each can help obtain the benefits of LfV (see Section 3.3). (3) Develop improved, novel action representations. Perhaps by combining the benefits of existing options.
@@
Previous LfV works have attempted to address LfV distribution shifts using specific inductive biases (see Section 4.2.2). However, these have not scaled to heterogenous internet video data. We advocate for implicitly addressing distribution shifts by scaling flexible methods to large, diverse internet video data, and, when possible, additionally training on (real and simulated) robot data. Compositional approaches are another option for obtaining generalization under distribution shift (?). Nevertheless, addressing LfV distribution shift in is an open problem.
@@
The high computational demands of training on video data can be mitigated via improved efficient architectures, taking inspiration from relevant LLM research (?, ?, ?, ?). The ever decreasing cost of compute will also help here (?). Improved methods for learning latent spaces in which the video model can operate (versus operating directly in pixel space), can aid computational efficiency in addition to mitigating issues related to noise and redundancy in video (?, ?).
6.4 Other Generalist Robot Challenges: Is Scaling Enough?
Observing results in other machine learning domains (?, ?, ?, ?), it seems likely that scaling current robot learning approaches to larger datasets will yield improvements in robot capabilities. This survey advocates for the use of internet video data to help us achieve this scaling. However, scaling may be insufficient to take us all the way to generalist robots (?). In this section, we note challenges that may face LfV approaches that rely on scaling the data under the current paradigm. Instead, solving these challenges may require algorithmic advances.
  • LfV is also reliant on robot data. LfV approaches augment a robot dataset with (internet) video data. However, any LfV approach will still likely be bottlenecked by the quantity, quality, and coverage of its robot data. The difficulties of collecting robot data may be the main barrier to matching the successes of deep learning in other domains, even for LfV approaches.
  • Safety and reliability. Safety is paramount when deploying robots in the real world. As such, physical robots may require higher success rates than required in many natural language tasks. Unfortunately, deep learning models are known to generalise poorly to unseen scenarios. Scaling to internet data may provide the model with improved ‘common-sense’, improving its reliability and generalization. However, this may not fundamentally solve the problem.
  • Long-horizon tasks and memory. Generalist robots may need to operate over long time horizons. This presents two challenges. First, longer-horizon tasks may necessitate improved planning and reasoning abilities. Second, longer-horizon tasks require improved memory capabilities. Current solutions to long-term memory involve scaling up the context-length of the model (?), or storing and retrieving from simple memory databases (?). It is unclear whether these solutions will be sufficient to fundamentally address the problem.
@@ 7 Conclusion
Developing general-purpose robots is a grand challenge in robotics. However, current learning approaches are bottlenecked by a lack of robot data. Learning from Video (LfV) methods seek to alleviate this issue by augmenting their training dataset with video data. These methods are promising as internet video comes in vast quantities and contains information highly relevant a general-purpose robot.
In this survey, we have outlined fundamental LfV concepts and conducted a comprehensive review of the LfV literature. We emphasised methods with the potential to scale to large, heterogeneous internet video datasets: following the success of scaling deep learning to internet data in other domains (?, ?), we consider these LfV methods to be highly promising for improving the generality of our robots.
@@
  • Fundamental LfV concepts. We explicitly establish the benefits that can be obtained from LfV (Section 3.3), and key LfV challenges that stand in the way of these benefits (Section 3.4). These fundamental concepts should be used to guide future LfV research agendas.
@@
  • Scalable approaches for learning policies and dynamics models. Our main analysis of the LfV-for-robotics literature (Section 4.3) yielded several important takeaways. First, future LfV research should avoid inductive biases that limit scalability (i.e., those seen in Section 4.2.2), and adopt scalable learning techniques similar to those used for video foundation models (Section 4.1). Whilst a reliance on language captions may limit scalability to an extent, this may become less of an issue as automated captioning methods improve (Section 5.1.2). Second, we should focus on methods that can best obtain the key LfV benefits (e.g., generalisation beyond the available robot data). This should involve targeting the learning of policies (Section 4.3.1) and dynamics models (Section 4.3.2) from video data. Improved benchmarks that quantitatively measure LfV benefits (see Section 5.2.2) will facilitate these efforts.
@@
  • Is scaling enough? Exploiting internet video will likely drive significant advancements in robotics. However, the generalist robot setting presents challenges that may not be solved via naive scaling of deep learning (Section 6.4). As LfV and other research attempts to address the data bottleneck in robotics, other complementary and alternative research agendas should be explored in parallel.

### v4

@@ 6.3 Overcoming LfV Challenges
The missing action label problem in LfV can be partially addressed using alternative-action representations (see Section 4.2.1). We recommend future research to: (1) Scale existing methods to realistic and diverse internet video. (2) Explicitly compare the pros and cons of different action representations, and measure the extent to which each can help obtain the benefits of LfV (see Section 3.3). (3) Develop improved, novel action representations. Perhaps by combining the benefits of existing options (?).
@@
Previous LfV works have attempted to address LfV distribution shifts using specific inductive biases (see Section 4.2.2). However, these have not scaled to heterogenous internet video data. We advocate for implicitly addressing distribution shifts by scaling flexible methods to large, diverse internet video data, and, when possible, additionally training on (real and simulated) robot data. Compositional approaches are another option for obtaining generalization under distribution shift (?). Nevertheless, addressing LfV distribution shift is an open problem.
@@
The high computational demands of training on video data can be mitigated via improved efficient architectures, taking inspiration from relevant LLM research (?, ?, ?, ?). The ever decreasing cost of compute will also help here (?). Improved methods for learning latent spaces in which the video model can operate (versus operating directly in pixel space) can aid computational efficiency in addition to mitigating issues related to noise and redundancy in video (?, ?).
6.4 Is Scaling Enough?
Observing results in other machine learning domains (?, ?, ?, ?), it seems likely that scaling current robot learning approaches to larger datasets will yield improvements in robot capabilities. This survey advocates for the use of internet video data to help us achieve this scaling. However, scaling may be insufficient to take us all the way to generalist robots (?). Indeed, there may be unique challenges in robotics that do not exist in other domains where scaling has proven successful (?). In this section, we note challenges that may face LfV approaches that rely on scaling the data under the current paradigm. Instead, solving these challenges may require algorithmic advances.
  • LfV is also reliant on robot data. LfV approaches augment a robot dataset with (internet) video data. However, any LfV approach will still be bottlenecked by the quantity, quality, and coverage of its robot data. The difficulty of collecting robot data with sufficient coverage over generalist robot settings may be the main barrier to scaling up robotics as seen in other domains, even for LfV approaches.
  • Safety and reliability. Safety is paramount when deploying robots in the real world. As such, physical robots may require higher success rates than are required in other generative AI tasks. Unfortunately, deep learning models are known to be brittle and to often generalise poorly to unseen scenarios (?, ?). Scaling to internet data may provide a robot model with improved ‘common-sense’, improving its reliability and generalization. However, this may not fundamentally solve the problem.
  • Long-horizon tasks and memory. Generalist robots may need to operate over long time horizons. This presents two challenges. First, longer-horizon tasks may necessitate improved planning, reasoning, and subtask success rates. Second, longer-horizon tasks require improved memory capabilities. Current solutions to long-term memory involve scaling up the context-length of the model (?), or storing and retrieving from simple memory databases (?). It is unclear whether these solutions will be sufficient to fundamentally address the problem.
@@ 6.4 Is Scaling Enough?
Nevertheless, ‘Is scaling enough?’ is ultimately an empirical question. Given deep learning will likely be a core component of generalist robot research for the foreseeable future, there is value in work pushing scaling-based approaches as far as possible. This will allow us to empirically answer this question and find fundamental and practical limitations that do indeed exist. Identified limitations can be tackled in a targeted manner by combining deep learning with alternative approaches (Section 2).
@@ 7 Conclusion
Developing general-purpose robots is a grand challenge in robotics. However, current robot learning approaches are bottlenecked by a lack of robot data. Learning from Video (LfV) methods seek to alleviate this issue by augmenting their training dataset with video data. These methods are promising as internet video comes in vast quantities and contains information highly relevant a general-purpose robot.
In this survey, we have outlined fundamental LfV concepts and conducted a comprehensive review of the LfV literature. We emphasised methods with the potential to scale to large, heterogeneous internet video datasets: following the success of scaling deep learning to internet data in other domains (?, ?), these LfV methods are promising for improving the generality of our robots.
@@
  • Fundamental LfV concepts. We explicitly establish the benefits that can be obtained from LfV (Section 3.3), and key LfV challenges that stand in the way of these benefits (Section 3.4). These fundamental concepts should guide future LfV research agendas.
@@
  • Scalable approaches for learning policies and dynamics models. Our main analysis of the LfV-for-robotics literature (Section 4.3) yielded important takeaways. First, future LfV research should avoid inductive biases that limit scalability (i.e., those seen in Section 4.2.2), and adopt scalable learning techniques similar to those used for video foundation models (Section 4.1). Though a reliance on language captions may limit scalability to an extent, this may become less of an issue as automated captioning methods improve (Section 5.1.2). Second, we should focus on methods that can best obtain the key LfV benefits (e.g., generalisation beyond the available robot data). This includes targeting the learning of policies (Section 4.3.1) and dynamics models (Section 4.3.2) from video data. Improved benchmarks that quantitatively measure LfV benefits (see Section 5.2.2) will facilitate these efforts.
@@
  • Is scaling enough? Exploiting internet video will likely drive significant advances in robotics. However, the generalist robot setting presents challenges that may not be solved via naive scaling of deep learning (Section 6.4). As LfV and other research attempts to address the data bottleneck in robotics, other complementary and alternative research agendas should be explored in parallel (see Section 2).